Grade 12Algebra

Give first the step you will use to separate the variable and then solve the equation:
x – 1 = 5

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

We have to add 1 to both the side of given equation,
Then we get,
= x - 1 +1 = 5 + 1
= x =6

Thanks